Bare Demo of IEEEtran.cls for Journals

Michael Shell, Member, IEEE, John Doe, Fellow, OSA, and Jane Doe, Life Fellow, IEEE

Abstract

The abstract goes here.

Index Terms

IEEEtran, journal, L^{A}TEX, paper, template.

I. Introduction

This demo file is intended to serve as a “starter file” for IEEE journal papers produced under L^{A}TEX using IEEEtran.cls version 1.6 and later. May all your publication endeavors be successful.

mds

August 13, 2002

A. Subsection Heading Here Subsection text here.

1) Subsubsection Heading Here: Subsubsection text here.

II. Conclusion

The conclusion goes here.

Appendix A

Proof of the First Zonklar Equation

Appendix one text goes here.

Manuscript received January 20, 2002; revised August 13, 2002. M. Shell is with the Georgia Institute of Technology.

August 29, 2008 DRAFT

Appendix B

Appendix two text goes here.

Acknowledgment

The authors would like to thank... This work was supported by the IEEE.

References

[1] H. Kopka and P.W. Daly, A Guide to L^{A}TEX, third ed. Harlow, U.K.: Addison-Wesley, 1999.

D.S. Coming and O.G. Staadt, ”Velocity-Aligned Discrete Oriented Polytopes for Dynamic Collision Detection,” IEEE Trans. Visualization and Computer Graphics, vol.14, no.1, pp. 1-12, Jan/Feb 2008, doi:10.1109/TVCG.2007.70405. H. Goto, Y. Hasegawa, and M. Tanaka, ”Efficient Scheduling Focusing on the Duality of MPL Representation,” Proc. IEEE Symp. Computational Intelligence in Scheduling (SCIS ’07), pp. 57-64, Apr. 2007, doi:10.1109/SCIS.2007.367670. J.M.P. Martinez, R.B. Llavori, M.J.A. Cabo, and T.B. Pedersen, ”Integrating Data Warehouses with Web Data: A Survey,” IEEE Trans. Knowledge and Data Eng., preprint, 21 Dec. 2007, doi:10.1109/TKDE.2007.190746.

PLACE PHOTO HERE Michael Shell Biography text here.

John Doe Biography text here.

Jane Doe Biography text here.

August 29, 2008 DRAFT